Brief Exercise 1-9 Your answer is partially correct. Try again. At the beginning of the year, Skysong, Inc. had total assets of $845,000 and total liabilities of $532,000. (Treat each item independently.) (a) If total assets increased $150,000 during the year and total liabilities decreased $75,000, what is the amount of stockholders equity at the end of the year? Stockholders equity 538000 (b) During the year, total liabilities increased $96,000 and stockholders equity decreased $80,000. What is the amount of total assets at the end of the year? Total assets (c) If total assets decreased $82,000 and stockholders equity increased $120,000 during the year, what is the amount of total liabilities at the end of the year? Total liabilities Click if you would like to Show Work for this question: Open Show Work
Answer #1
Assets = Liabilities + Equity
Given 845,000 = 532,000 + 313,000
a) 150,000 = -75,000    + 0
995,000 457,000 538,000
Stockholders Equity 538,000
b) Assets = Liabilities + Equity
Given 845,000 = 532,000 + 313,000
0 = 96,000    + -80,000
861,000 = 628,000 233,000
total assets= 861,000
c)
Assets = Liabilities + Equity
Given 845,000 = 532,000 + 313,000
-82,000 = 0    + 120,000
763,000 = 330,000 433,000
total liabilities= 330,000